Consider the following forwarding table.

Destination Address Range

Link Interface

10000011 00000001 00110000 00000000

through                                

10000011 00000001 00110111 11111111

0

10000011 00000001 00111000 00000000

through

10000011 00000001 00111111 11111111

1

10000011 00000001 00110000 00000000

through

10000011 00000001 00110001 11111111

2

10000011 00000001 00110010 00000000

through

10000011 00000001 00110011 11111111

3

Otherwise

4

a) For each of the following IP address, what does the router do if packets with the following addresses arrive? Explain why briefly.

  1. 131.1.64.1    
  2. 131.1.51.20   
  3. 131.1.49.7    
  4. 131.1.60.120  
  5. 131.1.52.59  
  6. b) Rewrite this forwarding table using the a.b.c.d/x notation instead of the binary string notation. Also determine the number of IP addresses for each network.

Destination Network Address (a.b.c.d/x)

Number of IP addresses

Link Interface

0

1

2

3

Otherwise

N/A

4
Answer #1

The given forwarding table has been converted to normal decimal form for better understanding. Please comment for further assistance.

Address Range Link Interface
131.1.48.0 to 131.1.55.255 0
131.1.56.0 to 131.1.63.255 1
131.1.48.0 to 131.1.49.255 2
131.1.50.0 to 131.1.51.255 3
Otherwise 4


a)

  1. 131.1.64.1 : This IP doesn't belongs to any of the given range. So, it forwards to interface 4.
  2. 131.1.51.20 : This IP can be valid with two entries with interfaces 0 and 3. But the range( subnet)  with less number of ip address will be selected to forward. So, here interface 3's range is less and it will be selected to forward.
  3. 131.1.49.7 : Like the previous question , this IP can be valid with two entries with interfaces 0 and 2. But, interface 2 will be selected because of it's size of range is small.
  4. 131.1.60.120 : This IP exclusively belongs to the range with interface 1 and forwarded accordingly.
  5. 131.1.52.59 : This IP address belongs to the range with interface 0 and forwarded accordingly.

b)

  

Destination Network Address Number of IP addresses Link interface
131.1.48.0 / 21 2^(32-21) = 2^11 = 2028 0
131.1.56.0 / 21 2^(32-21) = 2^11 = 2048 1
131.1.48.0 / 23 2^(32-23) = 2^9 = 512 2
131.1.50.0 / 23 2^(32-23) = 2^9 = 512 3
Othewise N/A 4